I would like your help to learn how to work out how to varify solutions to linear equations. I have no clue where to begin. I also don't understand how to graph systems of equations. Here's one of my problems. graph: x+2y= -10 ; -x+3y= -9. Thank you for your help.
Answer by checkley71(8403)

X+2Y=-10
-X+3Y=-9 ADDING THE TWO EQUATIONS WE GET
--------
5Y=-19 OR Y=-19/5 SUBSTITUTING -19/5 FOR Y IN EQUATION 1 WE GET
X+2(-19/5)=-10 OR X-38/5=-10 OR X=-10+38/5 OR X=(-50+38)/5 OR X=-12/5
PROOF -(-12/5)+3(-19/5)=-9 OR 12/5-57/5=-9 OR (12-57)/5=-9 OR -45/5=-9 OR
-9=-9 & -12/5+2(-19/5)=-10 or -12/5-38/5=-10 or (-12-38)/5=-10 or -50/5=-10 or
-10=-10